# Extending Douglas–Rachford Splitting for Convex Optimization

### v1

@@ 6.1 Extending ADMM
To arrive at an extended version of ADMM, we follow the derivation and notation of [ryu2022large, Section 3.1]. We specialize to the finite-dimensional case $\mathcal{H}=\mathbb{R}^{p}$, though we note that, under mild assumptions, we also have weak convergence in the general real Hilbert setting. Let $f:\mathbb{R}^{p}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathbb{R}^{q}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex functions. Let $A\in\mathbb{R}^{n\times p},B\in\mathbb{R}^{n\times q}$, and $c\in\mathbb{R}^{n}$. Consider the following convex optimization problem:
@@
subject to $Ax+By=c.$
which can be reformulated as the following unconstrained problem
@@
for each $z\in\mathbb{R}^{n}$ and analogously for $B\rhd g$. The classical ADMM method alternatingly minimizes the augmented Lagrangian of ˜14
@@
$\left\lfloor\begin{aligned} x^{k+1}&\in\operatorname*{Argmin}_{x\in\mathbb{R}^{p}}L_{\alpha}\left(x,y^{k},u^{k}\right)\\ y^{k+1}&\in\operatorname*{Argmin}_{y\in\mathbb{R}^{q}}L_{\alpha}(x^{k+1},y,u^{k})\\ u^{k+1}&=u^{k}+\alpha(Ax^{k+1}+By^{k+1}-c)\end{aligned}\right$. (16)
This method can be derived by applying unrelaxed Douglas–Rachford splitting (i.e., with $\theta=1$) either to ˜15 or to the corresponding dual problem
@@
By instead applying Algorithm˜1 to ˜17, and invoking the convergence result in Theorem˜3.3, we obtain new convergent variants of ADMM. With relaxation parameter $\theta=1$, we arrive at Algorithm˜2, which extends the classical ADMM in ˜16 by allowing for different penalty parameters $\alpha$ and $\beta$ in the two augmented Lagrangians. From Theorem˜3.1, we conclude, under mild additional assumptions, that Algorithm˜2 converges if $0<\alpha<2\beta$.
@@
For a general relaxation parameter $\theta\in(0,\min\{2,2\beta/\alpha\})$ we instead arrive at the following relaxed version of extended ADMM, see Algorithm˜3.
Algorithm 3 Extended ADMM (general)
@@
For the derivation of the extended ADMM, we use the following identity (see [ryu2022large, Equation (2.6)]). Let $h:\mathbb{R}^{m}\rightarrow\mathbb{R}\cup\{+\infty\}$ be proper, closed and convex and let $C\in\mathbb{R}^{n\times m}$. If $\operatorname*{relint}\operatorname*{dom}h^{*}\cap\operatorname*{ran}(-C^{\top})\neq\emptyset$ then
@@
$\psi^{k+1}$ $=\psi^{k}+\theta\frac{\alpha}{\beta}(\mu^{k+1}-\mu^{k+1/2}).$
We have that $J_{\alpha\partial\tilde{g}}(\cdot)=J_{\alpha\partial(g^{*}\circ(-B^{\top}))}(\cdot-\alpha c)$ by [BauschkeCombettes2017, Proposition 23.17]. Then by using ˜18 and defining auxiliary variables $\tilde{y}^{k+1}\in\mathbb{R}^{q}$ and $\tilde{x}^{k+1}\in\mathbb{R}^{p}$, the update of ˜19 becomes
@@
We can now remove $\mu^{k+1}$ and $\mu^{k+1/2}$ as they are redundant, make the substitution $u^{k}=\psi^{k}-\theta\alpha A\tilde{x}^{k}$, and reorganize the iterates to arrive at the update
@@
Remark 6.1:
If $(\alpha,\beta,\theta\alpha/\beta)\in S^{(1)}\cup S^{(2)}$ and the inclusion problem obtained from ˜17 is solvable and involves operators in $\partial\mathcal{F}\times\partial\mathcal{F}$ we have convergence by Theorem˜3.3. For the derivation, we furthermore require that the right implication of ˜18 holds both for $\tilde{f}$ and for $\tilde{g}$, that both functions are proper, closed and convex and that the minimization problems involved are solvable (see, for example, [ryu2022large] and [Chen_2016] for more discussion on regularity assumptions).
@@ 6.2 Extending Chambolle–Pock
To arrive at an extended version of the Chambolle–Pock method, we follow the derivation and notation of [o2020equivalence]. We specialize to the finite-dimensional case $\mathcal{H}=\mathbb{R}^{n}$, though we note that, under mild assumptions, we also have weak convergence in the general real Hilbert setting. Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex functions. Let $A\in\mathbb{R}^{m\times n}$. Consider the optimization problem
@@
In [he2012convergence, Algorithm 4], they propose the following relaxation for the Chambolle–Pock method with $\theta=1$:
@@
When the extended Douglas–Rachford method proposed in this paper is combined with the approach of [o2020equivalence], we obtain the doubly relaxed Chambolle–Pock method of Algorithm˜4, combining both types of relaxation. See Section˜6.2.1 for its derivation.
@@
1:Input: $(f,g,A)$ as in ˜20, $(x^{0},z^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$,
@@
This algorithm is guaranteed to converge on a strictly larger parameter region—namely $(\tau,\sigma,\theta,\rho)\in\mathbb{R}^{4}_{++}$ such that $\rho<\min\{2,2\theta\}$ and $\tau\sigma\left\lVert A\right\rVert^{2}\leq 1/\theta$—than that established in [o2020equivalence]. In fact, our analysis of Algorithm˜4 is tight in the sense that none of these two conditions can be relaxed without restricting the other one. To see this, let $n=m=1$, $A=1$, $f=g^{*}=0$. This defines an optimization problem of the form ˜20 with the unique minimizer 0, and gives the iteration
@@
where $v^{k}=(x^{k},z^{k})$. With the variable $\omega\coloneqq\tau\sigma\left\lVert A\right\rVert^{2}$, we let $\lambda_{-}$ denote the eigenvalue of ˜23 given by the expression
@@
This result reveals a trade-off between the two bounds: the conditions $\rho<\min\{2,2\theta\}$ and $\tau\sigma\left\lVert A\right\rVert^{2}\leq 1/\theta$ cannot be relaxed independently—making one upper bound larger necessarily forces the other to become more restrictive. However, if we, for instance, fix $\rho\in(0,\min\{2,2\theta\})$ to a specific value, the admissible step-size region can be enlarged. For example, for $\rho=1$ and $\theta>\tfrac{1}{2}$, the analysis in [banert2025chambolle] allows $\tau\sigma\left\lVert A\right\rVert^{2}<4/(1+2\theta)$, which is strictly less restrictive than our bound in this regime. It remains to be seen if the Chambolle–Pock method can be derived from the Douglas–Rachford method without conservatism in the step-size restriction for each individual choice of $\rho$.
@@
6.2.1 Derivation of Double-Relaxed Chambolle–Pock
@@
$0\in\partial\tilde{f}(u)+\tilde{G}(u),$ (25)
where $\tilde{G}\coloneqq B^{\top}\circ\partial g\circ B:\mathbb{R}^{n+m}\rightrightarrows\mathbb{R}^{n+m}$ is a maximally monotone operator.
Under mild assumptions, the monotone operator $\tilde{G}$ is the subdifferential of a proper, closed, and convex function. We will not need to assume that and instead apply Theorem˜3.3 to the pair $(\partial\tilde{f},\tilde{G})\in\partial\mathcal{F}\times\mathcal{A}$. Since
[…]
$\iff$ $(\exists u^{\star}_{1}\in\mathbb{R}^{n})\ \ \ \ \ u^{\star}_{1}\in\operatorname*{Argmin}(f+g\circ A),$
[…]
@@
$w^{k}$ $=J_{\eta\tilde{G}}\left(\left(1+\frac{\eta}{\tau}\right)\bar{u}^{k}-\frac{\eta}{\tau}y^{k}\right)$
@@
$\bar{w}^{k}$ $=J_{(\eta\tilde{G})^{-1}}\left(\left(1+\frac{\eta}{\tau}\right)\bar{u}^{k}-\frac{\eta}{\tau}y^{k}\right)$
@@
$J_{(\eta\tilde{G})^{-1}}(u)=\eta B^{\top}J_{\sigma\partial g^{*}}(\sigma Bu)$
@@ 6.3 Extending Parallel Splitting
$\operatorname*{minimize}_{y\in\mathcal{H}^{n}}f(y)+\iota_{\Delta}(y)$ (36)
with sufficient optimality condition
@@
7 Operator Composition Interpretation
In this section, we reinterpret our results through the lens of averaged, nonexpansive, and conic operators. We show that the algorithm operator defined by (4) need not be averaged when $\alpha\neq\beta$, and its convergence therefore lies outside the scope of classical averaged operator theory and the associated (preconditioned) proximal point framework. Readers not familiar with these operator-theoretic notions may safely skip this section, as it contains no new results.
We begin by recalling some basic definitions and properties of nonexpansive, averaged, and conic operators. An operator $N:\mathcal{H}\to\mathcal{H}$ is nonexpansive if $\|Nx-Ny\|\leq\|x-y\|$ for each $x,y\in\mathcal{H}$. For $\lambda\geq 0$, an operator $C:\mathcal{H}\to\mathcal{H}$ is called $\lambda$-conic if there exists a nonexpansive mapping $N:\mathcal{H}\to\mathcal{H}$ such that $C=(1-\lambda)\operatorname*{Id}+\lambda N$. If $\lambda\in[0,1)$ then $C$ is $\lambda$-averaged, if $\lambda=1$ then $C=N$ is nonexpansive, and if $\lambda>1$ then $C$ is not necessarily nonexpansive. Figure˜2 provides a graphical illustration of these classes. Given a mapping $R:\mathcal{H}\to\mathcal{H}$ and an iterate $z\in\mathcal{H}$, we say that the update $z^{+}:=(1-\mu)z+\mu R(z)$ with $\mu\in[0,1)$ is a $\mu$-averaged iteration of $R$. We also say that $T:=(1-\mu)\operatorname*{Id}+\mu R$ is a $\mu$-averaging of $R$. Since the iterates of an averaged operator converge weakly to a fixed point, this framework—equivalent to the relaxed proximal point framework [Martinet1970, Rockafellar1976PPA]—provides a powerful tool for analyzing fixed-point iterations, see, e.g., [DaoPhan2019AdaptiveDRS, GiselssonMoursi2021Compositions, BartzDaoPhan2022ConicalAveragedness, DavisYin2017ThreeOperator, BauschkeCombettes2017]. We also rely on the following result on compositions of conic operators.
Fact 7.1 ([GiselssonMoursi2021Compositions, Theorem 4.2, Proposition 4.5]):
Let $\lambda_{1},\lambda_{2}\geq 0$ and let $R_{1}:\mathcal{H}\to\mathcal{H}$ and $R_{2}:\mathcal{H}\to\mathcal{H}$ be $\lambda_{1}$-conic and $\lambda_{2}$-conic respectively.
  1. i)
    If $\lambda_{1}\lambda_{2}<1$, then $R_{2}R_{1}$ is $\frac{\lambda_{1}+\lambda_{2}-2\lambda_{1}\lambda_{2}}{1-\lambda_{1}\lambda_{2}}$-conic.
  2. ii)
    If $\max(\lambda_{1},\lambda_{2})=1$, then $R_{2}R_{1}$ is nonexpansive.
  3. iii)
    If $\lambda_{1}\lambda_{2}>1$, then there exists $\lambda_{1}$-conic $\widetilde{R}_{1}$ and $\lambda_{2}$-conic $\widetilde{R}_{2}$ such that $\widetilde{R}_{2}\widetilde{R}_{1}$ is not conic.
$zz^{\star}$$\lambda=\;$$\lambda=\;16383.99998$$\lambda=\;$16383.99998
Figure 2: Two-dimensional representations of three $\lambda$-conic operators. When $\lambda=0.5$, the operator is averaged (and therefore nonexpansive); when $\lambda=1.0$, the operator is nonexpansive but not averaged; and when $\lambda=1.5$, the operator is conic but not necessarily nonexpansive. For each operator $C:\mathbb{R}^{2}\to\mathbb{R}^{2}$ and any fixed point $z^{\star}\in\mathbb{R}^{2}$, evaluating $C$ at $z\in\mathbb{R}^{2}$ produces $C(z)\in\mathbb{R}^{2}$ lying within the corresponding shaded region. For averaged operators, the distance to all fixed points strictly decreases unless $C(z)=z$; for nonexpansive operators, the distance never increases; and for conic operators with $\lambda>1$, the distance may increase.
The algorithm in (3) with parameters $(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{2}\times(\mathbb{R}\setminus\{0\})$, repeated here for convenience
can equivalently be written as an averaged iteration of a composition of two conic operators. To see this, let us introduce the relaxed resolvents
$R_{A}^{\alpha,\beta}\coloneqq\left(1+\frac{\beta}{\alpha}\right)J_{\alpha A}-\frac{\beta}{\alpha}\operatorname*{Id}\qquad\quad{\text{and}}\qquad\quad R_{B}^{\beta,\alpha}\coloneqq\left(1+\frac{\alpha}{\beta}\right)J_{\beta B}-\frac{\alpha}{\beta}\operatorname*{Id},$ (39)
which allow us to rewrite the update in (38) as
$z^{+}=T_{A,B}^{\alpha,\beta,\theta}(z)\qquad\quad\text{where}\quad\qquad T_{A,B}^{\alpha,\beta,\theta}\coloneqq\left(1-\frac{\theta\beta}{\alpha+\beta}\right)\operatorname*{Id}+\frac{\theta\beta}{\alpha+\beta}R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}.$ (40)
Under the parameter restrictions in (5), we have $\frac{\theta\beta}{\alpha+\beta}\in[0,1)$ so (40) is a $\frac{\theta\beta}{\alpha+\beta}$-averaged iteration of the composition $R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}$. Each resolvent $J_{\alpha A}$ of a maximally monotone operator is firmly nonexpansive, equivalently $\tfrac{1}{2}$-averaged [BauschkeCombettes2017, Remark 4.34]. Therefore, for each $A\in\mathcal{A}$, there exists a nonexpansive $N:\mathcal{H}\to\mathcal{H}$ such that $J_{\alpha A}=\tfrac{1}{2}(\operatorname*{Id}+N)$, implying that

### v2

@@ 6.1 Extending ADMM
To arrive at an extended version of ADMM, we follow the derivation and notation of [ryu2022large, Section 3.1]. We specialize to the finite-dimensional case $\mathcal{H}=\mathbb{R}^{n}$, though we note that, under mild assumptions, we also have weak convergence in the general real Hilbert setting. Let $f:\mathbb{R}^{p}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathbb{R}^{q}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex functions. Let $A\in\mathbb{R}^{n\times p},B\in\mathbb{R}^{n\times q}$, and $c\in\mathbb{R}^{n}$. Consider the following convex optimization problem:
@@
subject to $Ax+By=c$
which can be reformulated as the following unconstrained problem,
@@
for each $z\in\mathbb{R}^{n}$ and analogously for $B\rhd g$. The classical ADMM method alternately minimizes the augmented Lagrangian of ˜15
@@
$\left\lfloor\begin{aligned} x^{k+1}&\in\operatorname*{Argmin}_{x\in\mathbb{R}^{p}}L_{\alpha}\left(x,y^{k},u^{k}\right)\\ y^{k+1}&\in\operatorname*{Argmin}_{y\in\mathbb{R}^{q}}L_{\alpha}(x^{k+1},y,u^{k})\\ u^{k+1}&=u^{k}+\alpha(Ax^{k+1}+By^{k+1}-c)\end{aligned}\right.$. (17)
This method can be derived by applying unrelaxed Douglas–Rachford splitting (i.e., with $\theta=1$) either to ˜16 or to the corresponding dual problem
@@
By instead applying Algorithm˜1 to ˜18, and invoking the convergence result in Theorem˜3.3, we obtain new convergent variants of ADMM. For a fixed relaxation parameter $\theta$, we arrive at Algorithm˜2, which extends the classical ADMM in ˜17 by allowing for different penalty parameters $\alpha$ and $\beta$ in the two augmented Lagrangians. From Theorem˜3.1, we conclude, under mild additional assumptions, that Algorithm˜2 converges if $0<\alpha<2\beta$. Using Algorithm˜2 it is also possible to derive new versions of proximal ADMM (see for example [ryu2022large] for a derivation with regular ADMM), and the preconditioned version of ADMM described in [BoydParikhChuPeleatoEckstein2011, Section 3.4.2], with different penalty parameters $\alpha$ and $\beta$.
@@
For a general relaxation parameter $\theta\in(0,\min\{2,2\beta/\alpha\})$ we instead arrive at the following extended version of the generalized ADMM [EcksteinBertsekas1992], see Algorithm˜3.
Algorithm 3 Extended Generalized ADMM
@@
For the derivation we use the following identity (see [ryu2022large, Equation (2.6)]). Let $h:\mathbb{R}^{m}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper, closed and convex function and let $C\in\mathbb{R}^{n\times m}$. If $\operatorname*{relint}\operatorname*{dom}h^{*}\cap\operatorname*{ran}(-C^{\top})\neq\emptyset$ then
@@
$\psi^{k+1}$ $=\psi^{k}+\frac{\theta\alpha}{\beta}(\mu^{k+1}-\mu^{k+1/2}).$
We have that $J_{\alpha\partial\tilde{g}}(\cdot)=J_{\alpha\partial(g^{*}\circ(-B^{\top}))}(\cdot-\alpha c)$ by [BauschkeCombettes2017, Proposition 23.17]. Then by using ˜19 and defining auxiliary variables $\tilde{y}^{k+1}\in\mathbb{R}^{q}$ and $\tilde{x}^{k+1}\in\mathbb{R}^{p}$, the update of ˜20 becomes
@@
We can now remove $\mu^{k+1}$ and $\mu^{k+1/2}$ as they are redundant, and make the substitution $u^{k}=\psi^{k}-\theta\alpha A\tilde{x}^{k}$, to arrive at the update
@@ 6.2 Extending Chambolle–Pock
To arrive at an extended version of the Chambolle–Pock method, we follow the derivation and notation of [o2020equivalence]. We specialize to the finite-dimensional case $\mathcal{H}=\mathbb{R}^{n}$, though we note that, under mild assumptions, we also have weak convergence in the general real Hilbert setting. Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex functions and let $A\in\mathbb{R}^{m\times n}$. Consider the optimization problem
@@
In [he2012convergence, Algorithm 4], they propose the following relaxation for the Chambolle–Pock method:
@@
When the extended Douglas–Rachford method proposed in this paper is combined with the approach of [o2020equivalence], we obtain the doubly relaxed Chambolle–Pock method of Algorithm˜4, combining both types of relaxations. See Section˜6.2.1 for its derivation.
@@
1:Input: $(f,g,A)$ as in ˜21, $(x^{0},z^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$,
@@
This algorithm is guaranteed to converge on a strictly larger parameter region—namely $(\tau,\sigma,\theta,\rho)\in\mathbb{R}^{4}_{++}$ such that $\rho<\min\{2,2\theta\}$ and $\tau\sigma\left\lVert A\right\rVert^{2}\leq 1/\theta$—than that established in [o2020equivalence]. In fact, our analysis of Algorithm˜4 is tight in the sense that neither of these two conditions can be relaxed without restricting the other. To see this, let $n=m=1$, $A=1$, $f=g^{*}=0$. This defines an optimization problem of the form ˜21 with the unique minimizer 0, and gives the iteration
@@
where $v^{k}=(x^{k},z^{k})$. With the variable $\omega\coloneqq\tau\sigma\left\lVert A\right\rVert^{2}$, we let $\lambda_{-}$ denote the eigenvalue of ˜24 given by the expression
@@
This result reveals a trade-off between the two bounds: the conditions $\rho<\min\{2,2\theta\}$ and $\tau\sigma\left\lVert A\right\rVert^{2}\leq 1/\theta$ cannot be relaxed independently—making one upper bound larger necessarily forces the other to become more restrictive. However, if we, for instance, fix $\rho\in(0,\min\{2,2\theta\})$ to a specific value, the admissible step-size region can be enlarged. For example, for $\rho=1$ and $\theta>1/2$, the analysis in [banert2025chambolle] allows $\tau\sigma\left\lVert A\right\rVert^{2}<4/(1+2\theta)$, which is strictly less restrictive than our bound in this regime. It remains to be seen if the Chambolle–Pock method can be derived from the Douglas–Rachford method without conservatism in the step-size restriction for each individual choice of $\rho$.
@@
6.2.1 Derivation of Doubly Relaxed Chambolle–Pock
@@
$0\in\partial\tilde{f}(u)+B^{\top}\circ\partial g\circ B(u),$ (26)
where, under mild assumptions, $\partial\tilde{g}=B^{\top}\circ\partial g\circ B:\mathbb{R}^{n+m}\rightrightarrows\mathbb{R}^{n+m}$ and solving the inclusion problem ˜26 is equivalent to solving the optimization problem ˜21, see [rockafellar1970convex, BauschkeCombettes2017].
@@
$w^{k}$ $={\rm{prox}}_{\eta\tilde{g}}\left(\left(1+\frac{\eta}{\tau}\right)\bar{u}^{k}-\frac{\eta}{\tau}y^{k}\right)$
@@
$\bar{w}^{k}$ $={\rm{prox}}_{(\eta\tilde{g})^{*}}\left(\left(1+\frac{\eta}{\tau}\right)\bar{u}^{k}-\frac{\eta}{\tau}y^{k}\right)$
@@
${\rm{prox}}_{(\eta\tilde{g})^{*}}(u)=\eta B^{\top}{\rm{prox}}_{\sigma g^{*}}(\sigma Bu)$
@@ 6.3 Extending Parallel Splitting
$\operatorname*{minimize}_{y\in\mathcal{H}^{n}}f(y)+\iota_{\Delta}(y),$ (36)
which is equivalent to (35), with sufficient optimality condition
@@
6.4 Extending Generalized Alternating Projections
In this section, we will derive an extended version of the generalized alternating projections method [GAP_Agmon, GAP_Motzkin, GAP_Bregman, falt-optimal, Dao_2018_Feasibility], which is also called the averaged relaxed alternating projections method or generalized Douglas–Rachford splitting.
Let $C\subset\mathcal{H}$ and $D\subset\mathcal{H}$ be nonempty, closed, and convex sets. We consider the problem
$\text{find }x\in\mathcal{H}\text{ such that }x\in C\cap D.$ (38)
Using the notation of ˜8, the Generalized Alternating Projections method has the update
$z^{k+1}=(1-\lambda)z^{k}+\lambda R_{D}^{\mu_{2}}R_{C}^{\mu_{1}}z^{k}$
for $\lambda\in(0,1]$ and $(\mu_{1},\mu_{2})\in(0,2)$ or $\lambda\in(0,1)$ and $(\mu_{1},\mu_{2})\in(0,2]$, see for example [falt-optimal]. In [Dao_2018_Feasibility] convergence is shown for larger values of the relaxation parameter $\lambda$.
By applying Algorithm˜1 to an appropriate reformulation, we get Algorithm˜6. This extends generalized alternating projections which is obtained by letting $\gamma=\beta/\alpha=1$ and requires $\mu_{1},\mu_{2}\in(0,2)$ and $\lambda\in(0,1)$. This is, to the best of the authors’ knowledge, the first projection method of this type that allows for relaxed projection steps that are longer than reflections.
Algorithm 6 Extended Generalized Alternating Projections
1:Input: Nonempty closed convex sets $C,D\subset\mathcal{H}$, $z^{0}\in\mathcal{H}$, and
[…]
3:for $k=0,1,2,\ldots$ do
4:  Update:
$z^{k+1}=(1-\lambda)z^{k}+\lambda R_{D}^{\mu_{2}}R_{C}^{\mu_{1}}z^{k}$
5:end for
The generalized alternating projections method can be derived by applying Douglas–Rachford splitting to the problem
$\operatorname*{minimize}_{x\in\mathcal{H}}\;c\operatorname{dist}_{C}^{2}(x)+d\operatorname{dist}_{D}^{2}(x),$ (39)
with constants $c,d\in(0,+\infty]$ and the interpretation that $(+\infty)\cdot\operatorname{dist}_{C}^{2}=\iota_{C}$. If $C\cap D$ is nonempty, this problem is equivalent to the feasibility problem ˜38, and the associated inclusion problem is solvable under mild assumptions (see for example [BauschkeCombettes2017]).
By applying Algorithm˜1 to ˜39, we instead get the Extended Generalized Alternating Projections algorithm stated in Algorithm˜6, which we will now derive.
The proximal operator of the squared distance function, with step size $\alpha\in\mathbb{R}_{++}$, satisfies
${\rm{prox}}_{\alpha c\operatorname{dist}_{C}^{2}}=\frac{c^{-1}}{c^{-1}+2\alpha}\operatorname*{Id}+\frac{2\alpha}{c^{-1}+2\alpha}\Pi_{C},$
and analogously for ${\rm{prox}}_{\beta d\operatorname{dist}_{D}^{2}}$ with the convention that $1/\infty=0$. The corresponding relaxed proximal operator defined in (8) becomes